OpenStudy (anonymous):

Car A has a positive velocity v(t) as it travels along a straight road, where v is a differentiable function of t. The velocity of the car is recorded for several selected values of t over the interval 0<=t<=60 seconds, as shown in the table below t (seconds): 0 10 20 30 40 50 60 v(t) (rate) : 5 14 7 11 12 40 44

OpenStudy (anonymous):

Car B travels along the same road with an acceleration of a(t) = 1/(t+9)^1/2 ft/sec2 At time t=0 seconds, the velocity of Car B is 3ft/sec. Which car is traveling faster at t=40 seconds? Background information that may or may not pertain to the question: Car A travels 1300 feet The acceleration at t=25 is 5/2

OpenStudy (anonymous):

I found the acceleration of Car A from 0 to 40, which was 40/7 I also found the acceleration of Car B from 0 to 40, which was -14 I got it by plugging in a(40), which was 1/7, for the equation and using that as my f(40) Am I on the right track? I got that Car A was faster but I don't think my process is right.

ganeshie8 (ganeshie8):

did u find the velocity of Car B ?

OpenStudy (anonymous):

No I didn't. Would I need to integrate 1/(t+9)^1/2 to do that?

ganeshie8 (ganeshie8):

use below :- velocity of Car B = \(\large \int a(t) dt\)

ganeshie8 (ganeshie8):

yess

ganeshie8 (ganeshie8):

and use the initial value (t=0, v=3) to find the constant

OpenStudy (anonymous):

Do I find the integral from 0 to 40?

ganeshie8 (ganeshie8):

we can find the indefinite integral, that gives u velocity function

OpenStudy (anonymous):

I got 2(t+9)^1/2 + C, but this is a definite integral If I find it from 0 to 40, I get 14-0 which is 14

OpenStudy (anonymous):

So if v(40) = 14, then . . . 40 - 0 / 14 - 3 = 40/11 which is actually smaller than 40/7, so Car A is faster. Does that look right to you?

ganeshie8 (ganeshie8):

wait a sec, definite integral between 0 and 40 gives u area under acceleration curve, wat does that represent ?

OpenStudy (anonymous):

Ah the distance I believe. What am I trying to find then?

ganeshie8 (ganeshie8):

you wanted to find velocity at t=40

OpenStudy (anonymous):

The velocity at t=40 was found from integrating the acceleration though right? Maybe the integral is not from 0 to 40, but when I plug 40 in for v(40) I get 14 right?

ganeshie8 (ganeshie8):

At time t=0 seconds, the velocity of Car B is 3ft/sec. Which car is traveling faster at t=40 seconds? \(\large v(t) = \int a(t) dt = \int 1/\sqrt{t+9} dt = 2 \sqrt{t+9} + c\) apply initial values : \(\large v(0) = 3\) \(\large 3 = 2\sqrt{0+9} + c \) \(\large c = -3\) velocity function : \(\large v(t) = 2 \sqrt{t+9} -3\)

ganeshie8 (ganeshie8):

plugin t = 40, to find velocity at t=40

OpenStudy (anonymous):

So v(40) = 11 Then 40 - 0 / 11 - 3 = 40/7, which is equal to car A.

ganeshie8 (ganeshie8):

nope, car A velocity at t=40 is 12

ganeshie8 (ganeshie8):

we simply pick the value from table directly

ganeshie8 (ganeshie8):

we are dealing wid 'instantaneous velocities'... not average...

OpenStudy (anonymous):

Okay let's see. OH that makes sense Okay well thank you very much. It was a little confusing at first but you've cleared it up. Thank you!

ganeshie8 (ganeshie8):

np :)